S. Kharibegashvili

A. Razmadze Mathematical Institute, Georgian Academy of Sciences Tbilisi, Georgia

ON SOME NONLOCAL PROBLEMS FOR A WAVE EQUATION

Some nonlocal problems are formulated and investigated for a wave equation in the case of two or three spatial variables.